# Augmented Object Intelligence: Making the Analog World Interactable with XR-Objects

### v4

@@ Abstract.
Seamless integration of physical objects as interactive digital entities remains a challenge for spatial computing. This paper introduces Augmented Object Intelligence  (AOI), a novel XR interaction paradigm designed to blur the lines between digital and physical by equipping real-world objects with the ability to interact as if they were digital, where every object has the potential to serve as a portal to vast digital functionalities. Our approach utilizes object segmentation and classification, combined with the power of Multimodal Large Language Models (MLLMs), to facilitate these interactions. We implement the AOI concept in the form of XR-Objects, an open-source prototype system that provides a platform for users to engage with their physical environment in rich and contextually relevant ways. This system enables analog objects to not only convey information but also to initiate digital actions, such as querying for details or executing tasks. Our contributions are threefold: (1) we define the AOI concept and detail its advantages over traditional AI assistants, (2) detail the XR-Objects  system’s open-source design and implementation, and (3) show its versatility through a variety of use cases and a user study.
@@
copyright: nonejournalyear: conference: ; ; booktitle: ccs: Human-centered computing Human computer interaction (HCI)ccs: Human-centered computing Mixed / augmented realityccs: Computing methodologies Intelligent agents
Figure 1. XR-Objects allows users to (a) select and interact with real-world objects in AR as if they were digital objects. (b) Automatically generated object-based AR context menus allow objects to provide information about themselves, such as nutritional facts. (c, d, e) For example, a user asks a question about cooking time, and (f, g) uses the result to set a spatial timer widget anchored to the relevant pot in 3D space.
@@
(a) A pack of pasta and a pot on a stove. (b) A menu panel overlaid on the pasta, which says: ”Filiz Burgu Pasta. (c) The interface says ”listening”. (d) ”how much time do I need to cook this pasta?” (e) Cook for 10 minutes” (f) The user taps ”timer” (g) A timer on the pot, anoter one on the teapot.
@@ 1. Introduction
In this paper, we explore a novel interaction paradigm we term Augmented Object Intelligence (AOI) that would allow any real object identified by the XR system to reveal digital data associated with the analog object and perform context-appropriate digital actions in a meaningful way. Our system XR-Objects  embodies this idea and aims to demonstrate and investigate “semantic equality” between real and virtual objects.
Imagine a scenario as familiar as right-clicking a digital file to open its context menu, but applied to physical objects within XR — such as right-clicking on potatoes or pasta in a pot to start a cooking timer set to the correct duration, or filtering for gluten-free products on a grocery shelf through an XR interface (see Figure 1).
The leap towards physical awareness in XR-Objects represents an advancement over traditional AR, which often relies on manual input or the use of physical tracking markers. We leverage developments in spatial understanding via technologies such as Simultaneous Localization and Mapping (SLAM), available in ARCore [^2] Google ARCore: https://developers.google.com/ar and ARKit [^3] Apple ARKit: https://developer.apple.com/augmented-reality/arkit, and machine learning models for object segmentation and classification (COCO (Lin et al., 2015) via MediaPipe (Lugaresi et al., 2019)). These technologies enable us to implement AR interactions with semantic depth. We also integrate a Multimodal Large Language Model (MLLM) into our system, which further enhances our ability to automate the recognition of objects and their specific semantic information within XR spaces.
@@
  • We introduce the concept of Augmented Object Intelligence (AOI), a paradigm shift towards seamless integration of real and virtual content in AIXR.
  • We detail the open-sourced design and implementation of XR-Objects, our prototypical system that exemplifies AOI, alongside an exploration of diverse use cases to demonstrate its potential.
  • We provide a comparison between standard prompt-based LLM interfaces and our AOI approach for contextual information retrieval and object-centric interaction.
We demonstrate our prototype that integrates these AR and AI components in a seamless way, which is implemented for smartphones to provide access to a broad audience, as commercial headsets do not yet give programmatic access to the user’s camera stream. By open-sourcing our system, we aim to foster further innovation in the field, ultimately bringing us closer to a future where the physical and digital realms interact seamlessly, embodying the concept of Programmable Reality (for anonymity during review process, 2024) for everyday users.
@@ 2. Related Work
This section provides an overview of previous work in the realms of user interface (UI) design principles, blended reality interactions, and advancements in AR technology (Speicher et al., 2019). These areas form the foundation upon which our research builds, which aims to enhance the interaction between users and the physical world through XR-Objects.
@@ 2.1. Fundamentals of UI
The challenge of bridging the cognitive gap between human users and computational systems has been a central theme in HCI. Traditional approaches have employed various layers of representations to mediate this interaction, manifesting most recognizably in the UIs of devices ranging from PCs and smartphones to IoT devices and automotive systems (Tidwell et al., 2020). Despite these advancements, the resurgence of command-line interface (CLIs) in contemporary AI interactions, as seen in the usage of prompts with large language models (LLMs) (Suh et al., 2024), suggests a potential oversimplification of user interaction paradigms. This regression shows the necessity of reevaluating our approach to UI design in the age of AI and spatial computing (Suzuki et al., 2023), where the intricacy of human intentions and the computational interpretation thereof demand a more nuanced form of representation. Figure 2 demonstrates that the multimodal AI assistant clearly has the capacity to produce reasonable scene understanding when an image and a prompt is provided as input, but it fails in providing an anchored output that ties back to the original multimodal prompt.
@@
2.2. Blended Reality
The exploration of blended reality systems is a rapidly developing field within HCI, marked by recent research that integrates real and virtual environments (Hartmann et al., 2019; Lindlbauer and Wilson, 2018; Lindlbauer et al., 2016). While a comprehensive literature review (Auda et al., 2023) exceeds the scope of this section, we draw attention to key themes that emerged within the broader XR literature.
@@ 2.2. Blended Reality
Balancing Immersion and External Awareness: Critical to blended-reality systems is managing the balance between immersion in the virtual world and awareness of the real world. Studies such as those by Kudo et al. emphasize the importance of smoothly transitioning users and devices between realities, while also maintaining bystander awareness (Kudo et al., 2021). Further guidelines on this balance are discussed in works by Gonzalez-Franco et al. for harmonizing user experience across dimensions (Gonzalez-Franco and Colaco, 2024; Gonzalez et al., 2024).
@@ 2.3. Interacting with Physical Objects in XR
Despite the considerable advances in AI’s capability to generate and understand complex content, our physical world remains predominantly analog, with only a fraction of daily activities and tools being enhanced by digital technology (Rogers et al., 2020). This analog nature of human experiences, from basic needs fulfillment to complex task execution, presents a significant challenge in integrating digital intelligence in a manner that feels both natural and easy to grasp.
Several research efforts have explored ways to bridge this gap by leveraging XR technologies to enable interaction with real-world objects. Figure 3 shows the landscape of physical object interactions in XR classified across two dimensions: anchoring (manual vs. seamless) and content (arbitrary vs. object-focused).
@@
Some approaches rely on pre-registration of objects or manual setups to achieve tangible input (Du et al., 2022; Monteiro et al., 2023; Zhu et al., 2022) or tangible haptic proxy (He et al., 2023; Jain et al., 2023; Hettiarachchi and Wigdor, 2016). In contrast to these manual processes, other works utilized markers to more automatically execute the object detection and AR content anchoring. For instance, researchers embedded visible (Rajaram and Nebeling, 2022; Henderson and Feiner, 2008; Chen et al., 2020; Li et al., 2019) or invisible markers (Dogan et al., 2022, 2023a, 2023b) to documents and 3D objects for AR purposes. However, these methods often impose limitations on the types of objects that can be interacted with, requiring specific fabrication or preparation.
@@
Once the objects have been identified and localized by the AR platform, an important dimension is what content will be shown at the identified locations (Caetano and Sra, 2022). Previous work has focused on optimizing the presentation of digital content (e.g., existing app windows or widgets) on or around objects (Cheng et al., 2021; Han et al., 2023; Lindlbauer et al., 2019), while others have investigated the use of tangible interaction with physical objects as a means to control such digital content (Zhu et al., 2022; Monteiro et al., 2023; Suzuki et al., 2020). EditAR (Chidambaram et al., 2022) suggested capturing users and their interactions with objects nearby to create digital twins for later consumption in XR. ProcessAR (Chidambaram et al., 2021) captures instruction demonstrations related to domain-specific objects by experts, so they can be viewed by novices in situ.
InfoLED (Yang and Landay, 2019), LightAnchors (Ahuja et al., 2019), and Reality Editor (Heun et al., 2013) took this one step further to show truly object-specific content in AR, however, this was specifically for electronic objects, such as their showing their battery level or device status in AR. Researchers further suggested augmenting visualizations with dynamic AR content (Chen et al., 2020; Chulpongsatorn et al., 2023a, b; Xiao et al., 2022), however, such dynamic content was limited to printed documents or displays.
@@ 3. XR-Objects Implementation
XR-Objects  leverages developments in spatial understanding via tools such as SLAM, available in ARCore [^4] Google ARCore: https://developers.google.com/ar and ARKit [^5] Apple ARKit: https://developer.apple.com/augmented-reality/arkit, and machine learning models for object segmentation and classification (COCO (Lin et al., 2015) via MediaPipe (Lugaresi et al., 2019)), which enables us to implement AR interactions with semantic depth. We also integrate a Multimodal Large Language Model (MLLM) into our system, which further enhances our ability to automate the recognition of objects and their specific semantic information within XR spaces.
@@ 3.1. Design Considerations
In contrast, our system prioritizes an object-centric approach, where interactions are directly anchored to objects within the user’s environment. This means that users can immediately access digital functionalities by selecting an object, without the need to navigate through an app or input additional information. By leveraging advanced computer vision and spatial understanding techniques, our AOI framework enables users to seamlessly engage with the physical world as if it were a digital interface.
The object-centric approach offers several advantages over app-centric models. Firstly, it provides a more natural interaction flow, as users can directly engage with objects in their surroundings without the cognitive burden of switching between the physical world and a digital app. Secondly, it minimizes the operational steps required to access digital functionalities, streamlining the user experience and reducing friction. Finally, by anchoring interactions directly to objects, our AOI framework fosters a more immersive and seamless XR experience.
@@ 3.2. Categories of Actions
In the above list, the first two categories (Information and Compare) represent traditional Visual Question Answering (VQA) tasks, while the other two (Share and Anchor) represent traditional widget tasks. We open-source our code on GitHub [^6] XR-Objects open-source project: https://github.com.anonymized and anticipate that the list of integrated actions will be extended in the future by the XR community.
@@ 3.3.1. Object Detection and Classification
Our algorithm also includes extra steps to ensure the object proxies do not get spawned at undesired locations or get erroneously duplicated for the same object.
@@ 3.3.4. Menu Interaction
4. Evaluation
We conducted a user study comparing XR-Objects to a state-of-the-art MLLM assistant interface (Gemini app [^9] Google Gemini: https://play.google.com/store/apps/details?id=com.google.android.apps.bard), referred to as ”Chatbot” from here on, for contextual information retrieval and object-centric interaction. Participants were asked to perform a number of timed VQA tasks and widget interactions in simulated grocery shopping and at-home scenarios, and provided feedback on their experience with each interface through a survey. This study was approved by Institutional Review Board at [anonymized].
4.1. Participants
We recruited 8 participants (6 male, 1 female, 1 preferred not to disclose) between the ages of 25-45. All were fluent English speakers (4 native), were regular shoppers, and all but 1 had used smartphone-based AR at least once.
4.2. Scenario & Tasks
We designed a scenario consisting of a simulated grocery shopping experience (Figure 6a) followed by an at-home experience (Figure 6b) in which users complete a set of 6 tasks using either XR-Objects  or Chatbot. The task categories included: (T1) 2-object Search, (T2) N-object Search, (T3) Share Object, (T4) Get Info (Single Object), (T5) Set/Anchor Timer, and (T6) Create/Anchor Note. We created two variants of this scenario (A and B), each with the same categories but slightly varied tasks. The full set of tasks is presented in Table 1.
@@ 4.2. Scenario & Tasks
For XR-Objects , participants were instructed to use the Compare feature for T1 and T2, the Share feature for T3, the Info feature for T4 and the Anchor feature for T5 and T6. For Chatbot, participants were instructed to take/upload a photo along with a query (using their preferred method of text or voice) to the chat for T1, T2, and T4. For the remaining tasks, participants were told to use Chatbot as if it were connected to a smartphone assistant.
4.3. Procedure
Participants were first given a brief introduction to the study, provided informed consent, and filled out a demographics survey. The experimenter then walked through the functionality of both XR-Objects  and Chatbot, and participants then completed a set of sample tasks on an object not included in the study.
@@ 4.3. Procedure
Each participant completed the tasks in both scenarios A and B. The ordering of both condition (XR-Objects,  Gemini) and scenario (AB) were counterbalanced between participants to prevent ordering effects. Once the tasks were completed, participants were free to use the tool freely for up to five additional minutes however they chose. Participants then completed a post-condition survey on their qualitative experience (Appendix A.1).
After a two minute break, this was repeated with the remaining condition and scenario. Upon completing the tasks with both conditions, participants completed a final survey comparing interactions with XR-Objects  and Chatbot (Appendix 4.4.3). Overall, the study took approximately 45 minutes.
4.4. Measures
4.4.1. Task Completion Time
We recorded the time required to complete tasks T1-T6 for each condition as a measure of overall system performance.
4.4.2. Post-Condition HALIE Survey
@@
4.4.3. Post-Study Form Factor Survey
While the user study was conducted on a smartphone due to limitations of head-mounted display (HMD) camera access, our vision for XR-Objects  is for it to run entirely on the HMD. Therefore, we conducted a post-study survey in which participants envisioned interacting with XR-Objects  on an HMD (e.g., Apple Vision Pro). The survey questions (provided in Appendix A.2) were based on HALIE, but formulated as a forced-choice comparison between the Chatbot and XR-Objects  interaction paradigms.
4.5. Analysis
4.5.1. Task Completion Time
We analyze completion time using traditional t-tests, and confirm normality via Shapiro-Wilk Test.
4.5.2. Distribution Skewness
Skewness ($\gamma_{1}$) quantifies the asymmetry of a given distribution’s shape. For a normal distribution, values of $|\gamma_{1}|$ 0.5 indicate an approximately symmetric distribution. Values of 0.5 ¡ $|\gamma_{1}|$ 1 suggest moderate skewness, while $|\gamma_{1}|$ 1 suggests a highly skewed distribution. This statistical approach, in contrast to visual methods like histograms, is particularly useful for analyzing distributions in Likert-scale questionnaires (King et al., 2018).
4.5.3. Generalized Linear Model

### v5

# Augmented Object Intelligence with XR-Objects

@@ Abstract.
Seamless integration of physical objects as interactive digital entities remains a challenge for spatial computing. This paper explores Augmented Object Intelligence  (AOI) in the context of XR, an interaction paradigm that aims to blur the lines between digital and physical by equipping real-world objects with the ability to interact as if they were digital, where every object has the potential to serve as a portal to digital functionalities. Our approach utilizes real-time object segmentation and classification, combined with the power of Multimodal Large Language Models (MLLMs), to facilitate these interactions without the need for object pre-registration. We implement the AOI concept in the form of XR-Objects, an open-source prototype system that provides a platform for users to engage with their physical environment in contextually relevant ways using object-based context menus. This system enables analog objects to not only convey information but also to initiate digital actions, such as querying for details or executing tasks. Our contributions are threefold: (1) we define the AOI concept and detail its advantages over traditional AI assistants, (2) detail the XR-Objects  system’s open-source design and implementation, and (3) show its versatility through various use cases and a user study.
@@
journalyear: 2024copyright: rightsretainedconference: The 37th Annual ACM Symposium on User Interface Software and Technology; October 13–16, 2024; Pittsburgh, PA, USAbooktitle: The 37th Annual ACM Symposium on User Interface Software and Technology (UIST ’24), October 13–16, 2024, Pittsburgh, PA, USAdoi: 10.1145/3654777.3676379isbn: 979-8-4007-0628-8/24/10ccs: Human-centered computing Human computer interaction (HCI)ccs: Human-centered computing Mixed / augmented realityccs: Computing methodologies Intelligent agents
Figure 1. XR-Objects  allows users to (a) select and interact with real-world objects in AR as if they were digital objects. Automatically generated object-based AR context menus allow objects to (b) provide information about themselves, such as nutritional facts and ingredients. For example, a user (c, d, e) asks a question about the cooking time of pasta, and then (f, g) uses the answer to set a spatial timer widget anchored to the relevant pot in 3D space.
@@
(a) A pack of pasta and a pot on a stove. (b) A menu panel overlaid on the pasta, which says: ”Filiz Burgu Pasta. (c) The interface says ”listening”. (d) ”how much time do I need to cook this pasta?” (e) Cook for 10 minutes” (f) The user taps ”timer” (g) A timer on the pot, another one on the teapot.
@@ 1. Introduction
In this paper, we explore an interaction paradigm we term Augmented Object Intelligence (AOI) in XR, which allows any analog object identified by the XR system to reveal digital data associated with it. This enables users to perform context-appropriate digital actions with respect to the object in a meaningful way. Our system XR-Objects  embodies this idea and aims to demonstrate and investigate “semantic equality” between real and virtual objects without the need for pre-registration.
Imagine a scenario as familiar as right-clicking a digital file to open its context menu, but applied to physical objects within XR — such as right-clicking on potatoes or pasta in a pot to start a cooking timer set to the correct duration, or filtering for gluten-free products on a grocery shelf through an XR interface (Figure 1).
The leap towards physical awareness in XR-Objects  represents an advancement over traditional AR, which often relies on manual input or the use of physical tracking markers. We uniquely combine developments in spatial understanding via technologies such as Simultaneous Localization and Mapping (SLAM), available in ARCore (Google, 2024a) and ARKit (Inc, 2024), and machine learning models for object segmentation and classification (COCO (Lin et al., 2015) via MediaPipe (Lugaresi et al., 2019)). These technologies enable us to implement object instance-based AR interactions with semantic depth and achieve live detection and 3D localization without pre-registration. We also integrate a Multimodal Large Language Model (MLLM) into our system, which further enhances our ability to automate the recognition of objects and their specific semantic information within XR spaces.
@@
  • We introduce the concept of Augmented Object Intelligence (AOI) for XR, a paradigm shift towards seamless integration of real and virtual content in XR using AI and object-based context menu interfaces.
  • We detail the open-sourced [^2] XR-Objects  open-source project: https://github.com/google/xr-objects design and implementation of XR-Objects, our prototypical system that exemplifies AOI, alongside an exploration of diverse use cases to demonstrate its potential.
  • We provide a comparative evaluation between standard prompt-based LLM interfaces and our AOI approach for contextual information retrieval and object-centric interaction in AR, highlighting AOI’s significant reduction in task completion time and its enhanced ease of use and satisfaction.
We demonstrate our prototype that integrates these AR and AI components in a seamless way, which is implemented for smartphones to provide access to a broad audience, as commercial XR headsets with large field of view (e.g., Meta Quest 3) do not yet give programmatic access to the user’s camera stream. By open-sourcing our system, we aim to foster further innovation in the field, ultimately bringing us closer to a future where the physical and digital realms interact seamlessly.
@@ 2. Related Work
This section provides an overview of previous work in user interface (UI) design principles, blended reality interactions, and advancements in AR technology (Speicher et al., 2019). These areas form the foundation upon which our research builds, which aims to enhance the interaction between users and the physical world through XR-Objects.
@@ 2.1. Fundamentals of UI
The challenge of bridging the cognitive gap between human users and computational systems has been a central theme in HCI. Traditional approaches have employed various layers of representations to mediate this interaction, manifesting most recognizably in the UIs of devices ranging from PCs and smartphones to IoT devices and automotive systems (Tidwell et al., 2020). Despite these advancements, the resurgence of command-line interface (CLIs) in contemporary AI interactions, as seen in the usage of prompts with large language models (LLMs) (Suh et al., 2024), suggests a potential oversimplification of user interaction paradigms. This regression shows the necessity of reevaluating our approach to UI design in the age of AI and spatial computing (Suzuki et al., 2023), where the intricacy of human intentions and the computational interpretation thereof demand a more nuanced form of representation. Figure 2 demonstrates that the multimodal AI assistant clearly has the capacity to produce reasonable scene understanding when an image and a prompt is provided as input, but it fails in providing a robustly anchored output or interaction capabilities that tie back to the original multimodal prompt.
@@
2.2. Extended Reality
XR systems represent a rapidly developing field within HCI, with recent works aiming to make the boundaries between real and virtual environments indistinguishable (Lindlbauer et al., 2016; Dogan et al., 2022a). While a comprehensive literature review (Auda et al., 2023) is beyond the scope of this section, we highlight key themes from the literature that inspired our research.
@@ 2.2. Extended Reality
Balancing Immersion and External Awareness: Critical to XR systems is managing the balance between immersion in the virtual world and awareness of the real world. Studies such as those by Kudo et al. emphasize the importance of smoothly transitioning users and devices between realities, while also maintaining bystander awareness (Kudo et al., 2021). Further guidelines on this balance are discussed in works by Gonzalez-Franco et al. for harmonizing user experience across dimensions (Gonzalez-Franco and Colaco, 2024; Gonzalez et al., 2024).
@@ 2.3. Interacting with Physical Objects in XR
Despite the considerable advances in AI’s capability to generate and understand complex content, our physical world remains predominantly analog, with only a fraction of daily activities and tools being enhanced by digital technology (Rogers et al., 2020; Wang and Zhang, 2024; Dogan et al., 2021). This analog nature of human experiences, from basic needs fulfillment to complex task execution, presents a significant challenge in integrating digital intelligence in a manner that feels both natural and easy to grasp.
Several research efforts have explored ways to bridge this gap by leveraging XR technologies to enable interaction with real-world objects (Xu et al., 2023). Figure 3 shows the landscape of physical object interactions in XR classified across two dimensions: anchoring (manual vs. seamless) and content (arbitrary vs. object-focused).
@@
Building on the foundational work in tangible bits (Ishii and Ullmer, 1997; Ishii, 2008), some approaches rely on pre-registration of objects or manual setups to achieve tangible input (Du et al., 2022; Monteiro et al., 2023; Zhu et al., 2022; Campos Zamora et al., 2024) or tangible haptic proxy (He et al., 2023; Jain et al., 2023; Hettiarachchi and Wigdor, 2016). In contrast to these manual processes, other works utilized markers to more automatically execute the object detection and AR content anchoring (Dogan, 2024). For instance, researchers embedded visible (Rajaram and Nebeling, 2022; Henderson and Feiner, 2008; Chen et al., 2020; Li et al., 2019) or invisible markers (Dogan et al., 2022b, 2023a, 2023b, 2020, c) to documents and 3D objects for AR purposes. However, these methods often impose limitations on the types of objects that can be interacted with, requiring specific fabrication or preparation.
@@
Once the objects have been identified and localized by the AR platform, an important dimension is what content will be shown at the identified locations (Caetano and Sra, 2022). Previous work has focused on optimizing the presentation of digital content (e.g., existing app windows or widgets) on or around objects (Cheng et al., 2021; Han et al., 2023; Lindlbauer et al., 2019; Lu and Xu, 2022), while others have investigated the use of tangible interaction with physical objects as a means to control such digital content (Zhu et al., 2022; Monteiro et al., 2023; Suzuki et al., 2020). EditAR (Chidambaram et al., 2022) suggested capturing users and their interactions with objects nearby to create digital twins for later consumption in XR. ProcessAR (Chidambaram et al., 2021) captures instruction demonstrations related to domain-specific objects by experts, so they can be viewed by novices in situ.
InfoLED (Yang and Landay, 2019), LightAnchors (Ahuja et al., 2019), BLEARVIS (Strecker et al., 2023), and Reality Editor (Heun et al., 2013) took this one step further to show truly object-specific content in AR, however, this was specifically for electronic objects, such as their showing their battery level or device status in AR. Researchers further suggested augmenting visualizations with dynamic AR content (Chen et al., 2020; Chulpongsatorn et al., 2023a, b; Xiao et al., 2022), however, such dynamic content was limited to printed documents or displays.
@@ 3. XR-Objects Implementation
XR-Objects  leverages developments in spatial understanding via tools such as SLAM, available in ARCore (Google, 2024a) and ARKit (Inc, 2024), and machine learning models for object segmentation and classification (COCO (Lin et al., 2015) via MediaPipe (Lugaresi et al., 2019)), which enables us to implement AR interactions with semantic depth. We also integrate a Multimodal Large Language Model (MLLM) into our system, which further enhances our ability to automate the recognition of objects and their specific semantic information within XR spaces.
@@ 3.1. Design Considerations
In contrast, our system prioritizes an object-centric approach, where interactions are directly anchored to objects within the user’s environment. This means that users can immediately access digital functionalities by selecting an object, without the need to navigate through an app or input additional information. By leveraging advanced computer vision and spatial understanding techniques, our AOI framework enables users to seamlessly engage with the physical world as if it were a digital interface. We also note that, currently, while our research prototype does need to be installed as an application package, it aims for eventual native integration, similar to how QR scanning is now embedded in smartphones. On XR headsets, AOI could be enabled as an additional layer of the home space, in a similar manner as video passthrough can be toggled on and off.
An object-centric approach would offer several advantages over app-centric models. Firstly, it provides a more natural interaction flow, as users can directly engage with objects in their surroundings without the cognitive burden of switching between the physical world and a digital app. Secondly, it minimizes the operational steps required to access digital functionalities, streamlining the user experience, enabling multi-tasking, and reducing friction. Finally, by anchoring interactions directly to objects, our AOI framework may offer a more immersive and seamless XR experience.
@@ 3.2. Categories of Actions
In the above list, the first two categories (Information and Compare) represent traditional Visual Question Answering (VQA) tasks, while the other two (Share and Anchor) represent traditional widget tasks. We open-source our code on GitHub [^3] XR-Objects  open-source project: https://github.com/google/xr-objects and anticipate that the list of integrated actions will be extended in the future by the XR community.
@@ 3.3.1. Object Detection and Classification
Our algorithm includes additional steps to prevent object proxies from being spawned in unintended locations or duplicated erro- neously for the same object.
@@ 3.3.4. Menu Interaction
4. Preliminary Evaluation
We conducted a user study comparing XR-Objects to a state-of-the-art MLLM assistant interface (Gemini app [^6] Google Gemini: https://play.google.com/store/apps/details?id=com.google.android.apps.bard), referred to as ”Chatbot” from here on, for contextual information retrieval and object-centric interaction. Participants were asked to perform a number of timed VQA tasks and widget interactions in simulated grocery shopping and at-home scenarios, and provided feedback on their experience with each interface through a survey.
4.1. Methods
Participants. We recruited 8 participants (6 male, 1 female, 1 preferred not to disclose) between the ages of 25-45. All were fluent English speakers (4 native), were regular shoppers, and all but 1 had used smartphone-based AR at least once.
Task & Procedure. We designed a scenario consisting of a simulated grocery shopping experience (Figure 6a) followed by an at-home experience (Figure 6b) in which users complete a set of 6 tasks using either XR-Objects  or Chatbot. The tasks (T1-T6) are listed in Table 1, and included retrieving/comparing information about multiple objects, sharing objects, and anchoring widgets.
@@ 4.1. Methods
For XR-Objects, participants were instructed to use the Compare feature for T1 and T2, the Share feature for T3, the Info feature for T4 and the Anchor feature for T5 and T6. For Chatbot, participants were instructed to take/upload a photo along with a query (using their preferred method of text and/or voice) to the chat for T1, T2, and T4. For the remaining tasks, participants were told to use Chatbot as if it were connected to a smartphone assistant.
First, participants were first given a brief introduction to the study, provided informed consent, and filled out a demographics survey. The experimenter then walked through the functionality of both XR-Objects  and Chatbot, and participants then completed a set of sample tasks on an object not included in the study.
@@
Each participant completed the tasks in both scenarios A and B. The ordering of both condition (XR-Objects,  Chatbot) and scenario (AB) were counterbalanced between participants to prevent ordering effects. Once the tasks were completed, participants were free to use the tool freely for up to five additional minutes however they chose. Participants then completed a post-condition survey on their qualitative experience.
After a two minute break, this was repeated with the remaining condition and scenario. Upon completing the tasks with both conditions, participants completed a final survey comparing interactions with XR-Objects  and Chatbot (Appendix LABEL:post-study_survey). Overall, the study took approximately 45 minutes.
Measures. As one of our measures, we recorded the time required to complete tasks T1-T6 for each condition as a measure of overall system performance.
@@
While the user study was conducted on a smartphone due to limitations of head-mounted display (HMD) camera access, our vision for XR-Objects  is for it to run entirely on the HMD. Therefore, we conducted a post-study form factor survey in which participants envisioned interacting with XR-Objects  on an HMD (e.g., Apple Vision Pro). The survey questions (provided in Appendix A.2) were based on HALIE, but formulated as a forced-choice comparison between the Chatbot and XR-Objects  interaction paradigms.
Analysis. We analyze completion time using traditional t-tests, and confirm normality via Shapiro-Wilk Test. Skewness ($\gamma_{1}$) quantifies the asymmetry of a given distribution’s shape. For a normal distribution, values of $|\gamma_{1}|$ 0.5 indicate an approximately symmetric distribution. Values of 0.5 ¡ $|\gamma_{1}|$ 1 suggest moderate skewness, while $|\gamma_{1}|$ 1 suggests a highly skewed distribution. This statistical approach, in contrast to visual methods like histograms, is particularly useful for analyzing distributions in Likert-scale questionnaires (King et al., 2018).